OpenStudy (anonymous):

find the missing terms of each arithmatic sequence 2,a2,a3,a4,-22,... please help!

jimthompson5910 (jim_thompson5910):

let a1 and a4 be the first and fourth terms respectively

jimthompson5910 (jim_thompson5910):

so a1 = 2 a2 = unknown a3 = unknown a4 = -22

jimthompson5910 (jim_thompson5910):

if d is the common difference, then to go from a1 to a2, you add on d like this a2 = a1 + d and to go from a2 to a3, you add on d again a3 = a2 + d and you repeat to get a4 a4 = a3 + d

jimthompson5910 (jim_thompson5910):

making sense so far?

OpenStudy (anonymous):

yea ok so i just need to do at as missing to and thn common sence shows the 4th term

OpenStudy (anonymous):

but im not sure how to get the common diffrence part

jimthompson5910 (jim_thompson5910):

so this means a2 = a1 + d a2 = 2 + d ... plug in a1 = 2 ------------- a3 = a2 + d a3 = (2+d) + d ... plug in a2 = 2+d a3 = 2+2d ------------- a4 = a3 + d a4 = (2+2d) + d ... plug in a3 = 2+2d a4 = 2 + 3d

jimthompson5910 (jim_thompson5910):

because a4 = 2 + 3d and because a4 = -22, we can say a4 = 2 + 3d -22 = 2 + 3d d = ??

OpenStudy (anonymous):

-8?

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

so a2 = a1 + d a2 = 2 + (-8) a2 = 2 - 8 a2 = -6

jimthompson5910 (jim_thompson5910):

what is a3?

OpenStudy (anonymous):

so i just add -8 to each term?

OpenStudy (anonymous):

-14?

jimthompson5910 (jim_thompson5910):

correct, and yes a3 = -14

OpenStudy (anonymous):

ok so i got 2,-6,-14,-22 but i still have a4 between -14 and -22?

jimthompson5910 (jim_thompson5910):

what do you mean when you say "but i still have a4 between -14 and -22"

OpenStudy (anonymous):

it says the correct answers are -4,-10,-16 i just need to show my work

jimthompson5910 (jim_thompson5910):

oh i messed up

OpenStudy (anonymous):

the question is: 2,a2,a3,a4,-22...

jimthompson5910 (jim_thompson5910):

I should have gone to a5, not a4

jimthompson5910 (jim_thompson5910):

solve for d a5 = a1 + 4d -22 = 2 + 4d d = ??

OpenStudy (anonymous):

-6

jimthompson5910 (jim_thompson5910):

so a2 = a1 + d a2 = 2 + (-6) a2 = 2 - 6 a2 = -4

OpenStudy (anonymous):

ok i got it thnx a bunch

jimthompson5910 (jim_thompson5910):

you're welcome